# DAWN-FM: Data-Aware and Noise-Informed Flow Matching for Solving Inverse Problems

### v2

@@ 3 Numerical Experiments
In this section, we experiment with our method on a few common datasets and two broadly applicable inverse problems: image deblurring and tomography (see A for details). We provide additional information on the experimental settings, hyperparameter choices, and network architectures in Tables 4, 5 and 3 for image deblurring task and Table 6 for tomography task in C.
@@
${\bf x}_{1}=\begin{bmatrix}{\bf x}_{1}^{\prime}\\ {\bf x}_{1}^{\prime}\end{bmatrix},\quad{\bf x}_{0}=\begin{bmatrix}{\bf x}_{0}^{\prime}\\ -{\bf x}_{0}^{\prime}\end{bmatrix}$ (19)
were generated by concatenation along the batch dimension. For a large number of samples, the sample mean of independent random variables converges to the true mean. However, the convergence can be slow due to the variance of the estimator being large. Antithetic sampling helped reduce this variance by generating pairs of negatively correlated samples, thereby improving convergence. The data was generated by employing the forward model along with a Gaussian noise injection, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$, where ${\bf z}\sim\mathcal{N}(0,{\bf I})$. The value of $\sigma$ was set to $p\%$ of the range of values in the data ${\bf b}$, where $p$ was sampled uniformly in (0, 20). ${\bf x}_{t}$ and ${\bf v}$ were computed from ${\bf x}_{1}$ and ${\bf x}_{0}$ following Equation 1 and Equation 2. Using the transformation $E_{\boldsymbol{\eta}}(f({\bf b}))=E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}$), the predicted velocity of the estimator was $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-FM estimator and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-FM estimator. Here, $E_{\boldsymbol{\eta}}$ was chosen to be a single-layer convolutional neural network. The loss for an epoch was computed as given in Equation 14.
@@
Baselines. For the image deblurring task, we compare our methods DAW-FM and DAWN-FM to diffusion-based image deblurring [Chung et al., 2022a] and InverseUNetODE [Eliasof et al., 2024]. The details for the training setup and hyperparameter choices for the baselines have been provided in C. All baselines were trained with Gaussian noise injection in the data in the same way as used for training of DAW-FM and DAWN-FM methods for fair comparison.
Metrics. For performance evaluation, we compute mean squared error (MSE), misfit, structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) between the ground truth and the reconstructed image. Additional details on these metrics are provided in B.
3.1 Image deblurring
We present the results of image deblurring task in Table 1 at noise level $p=5\%$ in blurred data for MNIST, STL10 and CIFAR10 datasets. For all metrics, DAWN-FM beats all other methods for all datasets by large margins, except for the CIFAR10 dataset, where DAW-FM performed marginally better than DAWN-FM. Moreover, on the CIFAR10 dataset, the diffusion model fits the data the best achieving the lowest value of the misfit metric. We also performed ablation studies for different noise levels in data for both DAW-FM and DAWN-FM models, as shown in Figure 5. For all metrics, DAWN-FM was more robust to noise in data, especially for higher noise levels over all datasets with the crossing point happening at values of $p\lesssim 5\%$ for all metrics. Some example images for the deblurring task for different levels of noise in data are presented in Figures 7, 8 and 9 in D. Based on empirical evidence from training, we believe that the results for both DAW-FM and DAWN-FM could be improved further with more training epochs.
Figure 5: Assessing the sensitivity of the metrics for the recovered deblurred image as a function of the level of noise ($p\%$) added to the data. The DAWN-FM is more robust to noise across all three datasets, MNIST, STL10 and CIFAR10 than DAW-FM, with the crossing point happening at values of $p$ $\lesssim$ $5\%$ for all metrics.
Table 1: Comparison between the performances of different methods (Diffusion, InverseUNetODE, and our methods DAW-FM and DAWN-FM) for the deblurring task for MNIST, STL10 and CIFAR10 datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for a large blurring kernel (for details, see A.1) at a noise level $p=5\%$ in the blurred data. Here, we report mean $\pm$ standard deviation over all images in the test set. First and second best performances are shown in bold and red, respectively.
@@ 3.1 Image deblurring
Figure 6: An example from OrganAMNIST dataset showing the different samples (labeled ${\bf x}^{(1)}_{1}$ to ${\bf x}^{(4)}_{1}$) drawn from the posterior starting from different initial ${\bf x}_{0}$ for both DAW-FM (top) and DAWN-FM (bottom). The posterior mean $\bar{{\bf x}}_{1}$ and standard deviation ${\sigma}_{{\bar{\bf x}}_{1}}$ were computed from 32 such samples, given the sinogram data (at noise level $p=5\%$). The red box highlights a lobe, which exhibits variability in its size and shape across different samples. This variability reflects the model’s uncertainty in reconstructing (the boundary of) such anatomical features.
3.2 Tomography
Table 2: Comparison between the performances of our methods (DAW-FM and DAWN-FM) for the tomography task for OrganAMNIST and OrganCMNIST datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for $N_{\text{angles}}=360$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the image dimension (for details, see A.2) at a noise level $p=5\%$ in the sinogram data. Here, we report mean $\pm$ standard deviation over 5000 images in the test set. Best performance is shown in bold.
@@ 3.2 Tomography
OrganAMNIST DAW-FM (Ours) 0.008 $\pm$ 0.002 0.041 $\pm$ 0.051 0.575 $\pm$ 0.108 20.389 $\pm$ 1.767
@@
OrganCMNIST DAW-FM (Ours) 0.014 $\pm$ 0.008 0.156 $\pm$ 1.424 0.511 $\pm$ 0.140 17.881 $\pm$ 2.624
@@
We present the results of our methods DAW-FM and DAWN-FM for the tomography task in Table 2 at noise level $p=5\%$ in the sinogram data for OrganAMNIST and OrganCMNIST datasets [Yang et al., 2023]. For this task, DAWN-FM outperformed DAW-FM on all metrics for both datasets. Some example images for the tomography task for different levels of noise are presented in Figures 10 and 11 in D. In medical image analysis, accurately estimating the size and boundaries of a lobe (such as the one shown in red box in Figure 6) is crucial for diagnostics, especially in cases where its size can influence medical decisions. We quantify uncertainty for both DAW-FM and DAWN-FM by computing mean and standard deviation across 32 samples from the learned posterior. While the mean represents the most probable reconstruction, the standard deviation map highlights regions of higher variability, indicating areas of uncertainty. From Figure 6, DAWN-FM gives a robust understanding of uncertainty by highlighting boundaries of objects as regions of highest uncertainty, a feature similar to what was observed on the MNIST dataset (see Figure 4).
4 Conclusions
@@
Appendix A Details about the Inverse Problems
A.1 Image deblurring
$K(x,y)=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\bigg{(}-\frac{x^{2}}{\sigma_{x}^{2}}-\frac{y^{2}}{\sigma_{y}^{2}}\bigg{)}$ (21)
The convolution operation can be described as a multiplication in the Fourier domain:
$B(u,v)=I(u,v)\cdot K(u,v)$ (22)
where $I(u,v)=\mathcal{F}\{I(x,y)\}$, $K(u,v)=\mathcal{F}\{K(x,y)\}$ are the Fourier transforms of the sharp image, and the PSF, respectively and $\mathcal{F}$ represents the Fourier transform. The blurred image can then be obtained using the inverse Fourier transform,
$I_{\text{blur}}(x,y)=\mathcal{F}^{-1}\{B(u,v)\}$ (23)
Moreover, ${\bf A}^{\top}$, the adjoint of operator ${\bf A}$, which is used to compute the data-embedding ${\bf A}^{\top}{\bf b}$ for the network, for some data ${\bf b}$ is easy to obtain. Since the image blurring operator is symmetric and self-adjoint, we have ${\bf A}={\bf A}^{\top}$.
A.2 Tomography
In tomography, the goal is to reconstruct the internal image of an object from a series of projections (sinograms) taken at various angles. This process can be described as solving an inverse problem, where the projections are obtained from the original image via the Radon transform ${\bf A}$ and inverse Radon transform ${\bf A}^{\top}$ is used to reconstruct the original image from the sinogram data.
The forward problem of tomography can be formulated as projecting a 2D image onto a set of 1D sinograms for different angles of view. The mathematical model of the forward projection is expressed as:
$S={\bf A}I_{\text{pad}}+\epsilon$ (24)
where, $S\in\mathbb{R}^{N_{\text{angles}}\times N_{\text{detectors}}}$ is the sinogram (the set of projections), $I_{\text{pad}}\in\mathbb{R}^{N_{\text{pixels}}}$ is the padded and flattened image of dimension $N_{\text{pixels}}$, ${\bf A}\in\mathbb{R}^{N_{\text{angles}}\times N_{\text{detectors}}\times N_{\text{pixels}}}$ is the tomography projection matrix which describes the interaction of each ray with every pixel, and $\epsilon$ is the additive noise term to the data, which was chosen as $p\%$ of the range of values in ${\bf A}I_{\text{pad}}$, with $p$ lying between 0% and 20% during training. For our experiments, we set $N_{\text{angles}}=360$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the dimension of the original unpadded images.
Similarly, the adjoint ${\bf A}^{\top}$ of the forward operator ${\bf A}$, can be used to recover the image, which can be expressed as,
$I_{\text{pad}}={\bf A}^{\top}S$ (25)
where, $I_{\text{pad}}\in\mathbb{R}^{N_{\text{pixels}}}$ is the (padded and flattened) reconstructed image, $S\in\mathbb{R}^{N_{\text{angles}}\times N_{\text{detectors}}}$ is the observed sinogram, ${\bf A}^{\top}\in\mathbb{R}^{N_{\text{pixels}}\times N_{\text{angles}}\times N_{\text{detectors}}}$ is the adjoint of the projection matrix. The $I_{\text{pad}}$ can be reshaped and unpadded to obtain an image of dimension $s\times s$.
Appendix B Evaluation Metrics
@@
B.1 Mean Squared Error (MSE)
The Mean Squared Error (MSE) measures the average squared difference between the pixel intensities of the original image and the reconstructed image. Given two images $I_{\text{true}}(x,y)$ (the ground truth image) and $I_{\text{rec}}(x,y)$ (the reconstructed image), MSE is calculated as:
$\text{MSE}=\frac{1}{H\cdot W}\sum_{x=1}^{H}\sum_{y=1}^{W}\left(I_{\text{true}}(x,y)-I_{\text{rec}}(x,y)\right)^{2}$ (26)
where:
@@ B.2 Misfit
$\text{Misfit}=\frac{1}{2}\sum_{x=1}^{H}\sum_{y=1}^{W}\left({\bf A}I_{\text{rec}}(x,y)-b(x,y)\right)^{2}$ (27)
where:
@@
B.3 Structural Similarity Index Measure (SSIM)
@@
$\text{SSIM}(I_{\text{true}},I_{\text{rec}})=\frac{(2\mu_{\text{true}}\mu_{\text{rec}}+C_{1})(2\sigma_{\text{true,rec}}+C_{2})}{(\mu_{\text{true}}^{2}+\mu_{\text{rec}}^{2}+C_{1})(\sigma_{\text{true}}^{2}+\sigma_{\text{rec}}^{2}+C_{2})}$ (28)
where:
@@
SSIM values range from $-1$ to $1$, where $1$ indicates perfect structural similarity.
B.4 Peak Signal-to-Noise Ratio (PSNR)
@@
$\text{PSNR}=-10\cdot\log_{10}(\text{MSE})$ (29)
wherel, MSE is the Mean Squared Error between the ground truth and reconstructed images. Higher PSNR values indicate better reconstruction quality.
Appendix C Experimental Settings
For the image deblurring task, the experiments were conducted on MNIST, STL10 and CIFAR10 datasets. The networks used were Diffusion model, InverseUNetODE as baselines and our proposed methods DAW-FM and DAWN-FM. The key details of the experimental setup for each of these methods are summarized in Tables 3, 4 and 5. For the tomography task, the experiments were conducted using DAW-FM and DAWN-FM methods on OrganAMNIST and OrganCMNIST datasets derived from the MedMNIST data library [Yang et al., 2023]. The key details of the experimental setup are summarized in Table 6. All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch [Paszke et al., 2017].
Table 3: Experimental details for the image deblurring task for our methods DAW-FM and DAWN-FM.

### v3

@@ 4. Numerical Experiments
In this section, we experiment with our method on a few common datasets and two broadly applicable inverse problems: image deblurring and tomography (see Section 3 for details). We provide additional information on the experimental settings, hyperparameter choices, and network architectures in Tables 5, 6 and 4 for image deblurring task and Table 7 for tomography task in Section 8.
@@
${\bf x}_{1}=\begin{bmatrix}{\bf x}_{1}^{\prime}\\ {\bf x}_{1}^{\prime}\end{bmatrix},\quad{\bf x}_{0}=\begin{bmatrix}{\bf x}_{0}^{\prime}\\ -{\bf x}_{0}^{\prime}\end{bmatrix}$, (26)
were generated by concatenation along the batch dimension. For a large number of samples, the sample mean of independent random variables converges to the true mean. However, the convergence can be slow due to the variance of the estimator being large. Antithetic sampling helped reduce this variance by generating pairs of negatively correlated samples, thereby improving convergence. The data was generated by employing the forward model along with a Gaussian noise injection, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$, where ${\bf z}\sim\mathcal{N}(0,{\bf I})$. The value of $\sigma$ was set to $p\%$ of the range of values in the data $\bf b$}, where $p$ was sampled uniformly in (0, 20). ${\bf x}_{t}$ and $\bf v$} were computed from ${\bf x}_{1}$ and ${\bf x}_{0}$ following Equation 1 and Equation 2.
Network implementation. To represent the temporal dynamics of flow matching and the effect of measurement noise within a unified network, we embed both the interpolation time $t$ and the noise level $\sigma$ into learnable feature vectors that modulate the network at every resolution. scalar time input $t\in[0,1]$ is passed through a linear projection to produce a time embedding, which is then broadcast and added as a per-channel bias at each of the encoder, bottleneck, and decoder blocks. This allows the network to adapt its feature transformations depending on the location on the flow trajectory. Similarly, the scalar noise level $\sigma$ is embedded into a noise feature vector. At each resolution, this vector is reshaped and combined multiplicatively with the backprojected data ${\bf A}^{\top}{\bf b}$, producing a noise-aware, spatially structured modulation of the features. In practice, this mechanism enables the model to rely more strongly on data consistency when the noise is low, while leaning more on the learned prior when the noise level is high. Together, the time embedding provides trajectory awareness and the noise embedding provides data-aware regularization, yielding a network that is both flow- and noise-adaptive.
To incorporate information from the measured data into the network at every resolution, we introduce a dedicated data embedding pathway that operates on the backprojected measurements ${\bf A}^{\top}{\bf b}$. At each encoder and decoder level, ${\bf A}^{\top}{\bf b}$ is first rescaled to match the current spatial resolution of the feature maps. This rescaled tensor is then processed by a lightweight convolutional module consisting of two consecutive $3\times 3$ convolutions with SiLU activations (overall denoted as $E_{\boldsymbol{\eta}}(\cdot)$ in Equation 15). This produces a set of feature maps that encode the structure of the backprojected measurements in a form that can be directly combined with the learned network features. By injecting these data features additively into each block, the network is explicitly encouraged to maintain coherence with the measured data throughout the forward and backward passes. In combination with the noise-conditioned modulation described above, the data embedding serves as the primary mechanism through which the measured observations influence the learned velocity field. Note that using the transformation $E_{\boldsymbol{\eta}}(f({\bf b}))=E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}$), the predicted velocity of the estimator was $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-FM estimator (no noise embedding) and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-FM estimator. The loss for an epoch was computed as given in Equation 14.
@@
Baselines. For the image deblurring task, we compare our methods DAW-FM and DAWN-FM to diffusion-based image deblurring [chung2022diffusion] and InverseUNetODE [eliasof2024completedeeplearningmethod]. The details for the training setup and hyperparameter choices for the baselines have been provided in Section 8. All baselines were trained with Gaussian noise injection in the data in the same way as used for training of DAW-FM and DAWN-FM methods for fair comparison.
Metrics. For performance evaluation, we compute mean squared error (MSE), misfit, structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) between the ground truth and the reconstructed image. Additional details on these metrics are provided in Section 6.
4.1. Image deblurring
We present the results of image deblurring task in Table 1 at noise level $p=5\%$ in blurred data for MNIST, STL10 and CIFAR10 datasets. For all metrics, DAWN-FM beats all other methods for all datasets by large margins, except for the CIFAR10 dataset, where DAW-FM performed marginally better than DAWN-FM. Moreover, on the CIFAR10 dataset, the diffusion model fits the data the best achieving the lowest value of the misfit metric. We also performed ablation studies for different noise levels in data for both DAW-FM and DAWN-FM models, as shown in Figure 6. For all metrics, DAWN-FM was more robust to noise in data, especially for higher noise levels over all datasets with the crossing point happening at values of $p\lesssim 5\%$ for all metrics. Some example images for the deblurring task for different levels of noise in data are presented in Figures 8, 9 and 10 in Section 9. Based on empirical evidence from training, we believe that the results for both DAW-FM and DAWN-FM could be improved further with more training epochs.
Figure 6. Assessing the sensitivity of the metrics for the recovered deblurred image as a function of the level of noise ($p\%$) added to the data. The DAWN-FM is more robust to noise across all three datasets, MNIST, STL10 and CIFAR10 than DAW-FM, with the crossing point happening at values of $p$ $\lesssim$ $5\%$ for all metrics.
Table 1. Comparison between the performances of different methods (Diffusion, InverseUNetODE, and our methods DAW-FM and DAWN-FM) for the deblurring task for MNIST, STL10 and CIFAR10 datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for a large blurring kernel (for details, see Section 3.1) at a noise level $p=5\%$ in the blurred data. Here, we report mean $\pm$ standard deviation over all images in the test set. First and second best performances are shown in bold and red, respectively.
@@ 4.1. Image deblurring
[…]
4.2. Tomography
Table 2. Comparison between the performances of our methods (DAW-FM and DAWN-FM) for the tomography task for OrganAMNIST and OrganCMNIST datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for $N_{\text{angles}}=180$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the image dimension (for details, see Section 3.2) at a noise level $p=5\%$ in the sinogram data. Here, we report mean $\pm$ standard deviation over 5000 images in the test set. First and second best performances are shown in bold and red, respectively.
@@ 4.2. Tomography
OrganAMNIST FBP (ramp filter) 0.014 $\pm$ 0.006 0.029 $\pm$ 0.006 0.469 $\pm$ 0.114 18.047 $\pm$ 2.129
DAW-FM (Ours) 0.008 $\pm$ 0.002 0.041 $\pm$ 0.051 0.575 $\pm$ 0.108 20.389 $\pm$ 1.767
@@
OrganCMNIST FBP (ramp filter) 0.015 $\pm$ 0.006 0.029 $\pm$ 0.008 0.437 $\pm$ 0.128 17.413 $\pm$ 2.271
DAW-FM (Ours) 0.014 $\pm$ 0.008 0.156 $\pm$ 1.424 0.511 $\pm$ 0.140 17.881 $\pm$ 2.624
@@
We present the results of our methods DAW-FM and DAWN-FM for the tomography task and compare them to FBP with ramp filtering in Table 2 at noise level $p=5\%$ in the sinogram data for OrganAMNIST and OrganCMNIST datasets [medmnistv2]. For this task, DAWN-FM outperformed DAW-FM on all metrics for both datasets. Some example images for the tomography task for different levels of noise are presented in Figures 11 and 12 in Section 9. In medical image analysis, accurately estimating the size and boundaries of a lobe (such as the one shown in red box in Figure 7) is crucial for diagnostics, especially in cases where its size can influence medical decisions. We quantify uncertainty for both DAW-FM and DAWN-FM by computing mean and standard deviation across 32 samples from the learned posterior. While the mean represents the most probable reconstruction, the standard deviation map highlights regions of higher variability, indicating areas of uncertainty. From Figure 7, DAWN-FM gives a robust understanding of uncertainty by highlighting boundaries of objects as regions of highest uncertainty, a feature similar to what was observed on the MNIST dataset (see Figure 5).
5. Conclusions
@@
Acknowledgments
S. Ahamed would like to acknowledge support from the Canadian Institutes of Health Research (CIHR) Grant PIBH-GR018169. Computational resources were provided by E. Haber.
Supplementary Material
6. Evaluation Metrics
@@
6.1. Mean Squared Error (MSE)
The Mean Squared Error (MSE) measures the average squared difference between the pixel intensities of the original image and the reconstructed image. Given two images $I_{\text{true}}(x,y)$ (the ground truth image) and $I_{\text{rec}}(x,y)$ (the reconstructed image), MSE is calculated as,
$\text{MSE}=\frac{1}{H\cdot W}\sum_{x=1}^{H}\sum_{y=1}^{W}\left(I_{\text{true}}(x,y)-I_{\text{rec}}(x,y)\right)^{2},$ (27)
where,
@@ 6.2. Misfit
$\text{Misfit}=\frac{1}{2}\sum_{x=1}^{H}\sum_{y=1}^{W}\left({\bf A}I_{\text{rec}}(x,y)-b(x,y)\right)^{2},$ (28)
where,
@@
6.3. Structural Similarity Index Measure (SSIM)
@@
$\text{SSIM}(I_{\text{true}},I_{\text{rec}})=\frac{(2\mu_{\text{true}}\mu_{\text{rec}}+C_{1})(2\sigma_{\text{true,rec}}+C_{2})}{(\mu_{\text{true}}^{2}+\mu_{\text{rec}}^{2}+C_{1})(\sigma_{\text{true}}^{2}+\sigma_{\text{rec}}^{2}+C_{2})},$ (29)
where,
@@
SSIM values range from 1-1 to $1$, where $1$ indicates perfect structural similarity.
6.4. Peak Signal-to-Noise Ratio (PSNR)
@@
$\text{PSNR}=-10\cdot\log_{10}(\text{MSE}),$ (30)
where, MSE is the Mean Squared Error between the ground truth and reconstructed images. Higher PSNR values indicate better reconstruction quality.
7. Additional Experiments with Inverse Crime Considerations
To examine whether our results are affected by the inverse crime [wirgin2004inverse], we conducted additional experiments in which the forward model used for data generation differed from that used during inversion. Specifically, instead of generating synthetic data directly using $\bf A$} as ${\bf b}={\bf A}{\bf x}_{1}+{\boldsymbol{\varepsilon}}$, we defined a composite forward operator,
$\tilde{{\bf A}}={\bf D}{\bf A}{\bf U},$ (31)
where $\bf U$} is an upsampling operator (by a factor of 2) applied to the clean image ${\bf x}_{1}$ and $\bf D$} is a downsampling operator (by a factor of 2) applied after the forward operation. This composite operator introduces a domain mismatch between the forward model used for data generation and the one used for inversion (which remains $\bf A$}), thereby eliminating the inverse crime. The results in Table 3 demonstrate that DAWN-FM maintains comparable reconstruction quality even under these more challenging conditions - for example, on the MNIST deblurring task, performance metrics such as MSE, SSIM, and PSNR remain nearly unchanged, and similar robustness is observed for the tomography task on OrganCMNIST. These results empirically confirm that the effectiveness of DAWN-FM seems to be robust to any implicit inverse crime, and the method generalizes well even when the forward model used for data generation is perturbed.
Table 3. Performance of DAWN-FM for the deblurring (on MNIST) and tomography (on OrganCMNIST) tasks with (using data generation forward problem, $\bf A$}) and without (using data generation forward problem, ${\bf D}{\bf A}{\bf U}$) inverse crime. Results show negligible degradation when a composite forward model slightly different from $\bf A$} is used for data generation.
Task (Dataset) Methods
Forward
model
Inverse crime present? MSE ($\downarrow$) MISFIT ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
Generation Inversion
Deblurring (MNIST) DAWN-FM $\bf A$} $\bf A$} Yes 0.073 $\pm$ 0.041 0.032 $\pm$ 0.009 0.901 $\pm$ 0.042 22.277 $\pm$ 2.612
${\bf D}{\bf A}{\bf U}$ $\bf A$} No 0.070 $\pm$ 0.034 0.041 $\pm$ 0.095 0.870 $\pm$ 0.090 22.198 $\pm$ 1.816
Tomography (OrganCMNIST) $\bf A$} $\bf A$} Yes 0.007 $\pm$ 0.007 0.151 $\pm$ 1.287 0.675 $\pm$ 0.116 21.630 $\pm$ 3.199
${\bf D}{\bf A}{\bf U}$ $\bf A$} No 0.009 $\pm$ 0.062 0.162 $\pm$ 1.113 0.632 $\pm$ 0.175 20.489 $\pm$ 3.795
8. Experimental Settings
For the image deblurring task, the experiments were conducted on MNIST, STL10 and CIFAR10 datasets. The networks used were Diffusion model, InverseUNetODE as baselines and our proposed methods DAW-FM and DAWN-FM. The key details of the experimental setup for each of these methods are summarized in Tables 4, 5 and 6. For the tomography task, the experiments were conducted using DAW-FM and DAWN-FM methods on OrganAMNIST and OrganCMNIST datasets derived from the MedMNIST data library [medmnistv2]. The key details of the experimental setup are summarized in Table 7. All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch [paszke2017automatic].
Table 4. Experimental details for the image deblurring task for our methods DAW-FM and DAWN-FM.